# Protecting Bystander Privacy via Selective Hearing in Audio LLMs

### v2

@@ 5.2 Evaluation
SH-Bench evaluation questions use a 5-way classification format, including an “I don’t know” option (or equivalent). With these questions, we evaluate the models under two different modes: i) General mode, where we instruct the model to answer the question given by the audio, and the model is expected to answer all questions correctly regardless of whether they concern the main speaker or the bystander; and ii) Selective mode, where we instruct the model that there is a main speaker in the foreground and provide a content-conditioned description of this speaker (see §3.2). The model is also instructed to only listen to the main speaker and ignore any background speech, so it should choose the IDK option for the questions related to what the bystander said, and should give correct answer for the main speaker. The prompts used in the evaluation are provided in Appendix A.1, and the evaluation procedure is shown with examples in Fig. 3.
Figure 3: Illustration of how accuracy is measured for the main speaker and the bystander in two modes. The main speaker must give the correct answer in both modes, whereas the bystander is only correct when the audio LLM gives the correct answer in general mode but selects “IDK” in the selective mode.
Therefore, accuracies measured for bystander under selective mode should treat the “I don’ t know" option as the correct choice. In addition to accuracies, we define Selective Efficacy as a unified metric using the harmonic mean of 4 different accuracies on the main speaker and bystanders under general or selective modes as follows.
$\text{SE}=\frac{4}{\sum_{m\in\{\text{gen., sel.}\}}\sum_{n\in\{\text{main, by.}\}}\text{Acc}_{\text{m,n}}^{-1}}$ (1)
This metric is high only when all accuracies are high, and there is no shortcoming. A high accuracy under general mode and a low accuracy under selective mode indicates the model has good audio understanding ability but poor privacy protection, whereas a low accuracy under general mode indicates the model being incompetent to comprehend audio content under multi-speaker and overlapped speech scenario. Besides, a high accuracy on bystander under selective mode with a low accuracy on main indicates the model being unable to distinguish main or bystander, and choose “I don’t know" regardless of the question.
@@ 6.1 Main Benchmark Results
Llama-Omni 2 14B fang2025llamaomni2 95.7 32.9 15.3 87.0 34.0
Qwen-2.5-Omni 7B qwen25 96.0 95.5 48.2 47.6 63.9
Qwen-2.5-Omni 3B qwen25 96.2 95.6 53.1 54.7 69.0
Step-Audio-2-mini stepaudio2 94.2 93.7 54.7 31.5 56.1
Kimi-Audio 7B Instruct kimiaudio 96.9 96.3 67.4 31.4 59.4
@@
Gemini 2.5 Pro gemini25 97.3 97.0 65.5 59.2 75.8
GPT-4o-audio-preview hurst2024gpt 72.3 84.4 43.2 44.5 56.1
@@
Step-Audio-2-mini + BPFT (ours.) 97.4 94.3 81.0 96.1 91.7
Qwen-2.5-Omni 7B + BPFT (ours.) 93.3 92.7 82.0 93.8 90.2
@@
Table 2: Model performances on SH-Bench test set under general mode (all questions should be answered correctly) and selective mode (bystander-privacy related questions should give I don’t know response). SE stands for Selective Efficacy which is the harmonic mean of all 4 accuracies. BPFT stands for bystander privacy fine-tuning introduced in Section 4. All metrics are the higher the better.
@@
The main results of models on SH-Bench are shown in Table 2, including the main and bystander speaker accuracies under general mode and selective mode respectively. The results using the BPFT stage are also highlighted.
@@
As the main speaker speech in our benchmark is relatively clean, the pronunciations are clear and the questions are direct, the accuracy for the main speaker is expected to be high for both general and selective mode. However, when it comes to bystander, due to smaller volume, speech overlap and background noise, it is more challenging to understand the speech content and hence the results are mixed. Gemini 2.5 Pro achieves the best performance on understanding bystander-related questions, with 65.5% accuracy under general mode.
@@
For selective mode, the task is mainly to test if the model can follow instructions and clearly distinguish which content is said by the main speaker and which is by the bystander. Since we provide “I don’t know" as one option, an over-reserved model may always choose “I don’t know" without actually understand the content (e.g. Llama-Omni-2) and often performs worse on main speakers as well under selective mode. Therefore, we report SE to reflect the selective hearing ability of all models by balancing all 4 accuracies. For existing models, Gemini 2.5 Pro achieved the best performance with SE of 75.8%, clearly higher than any other systems.
@@
Real vs. synthetic scenarios: Accuracies on real and synthetic scenarios are given in Table 6 in Appendix A.2. Since real scenarios are more challenging, we observed consistently lower accuracies on real scenarios for bystander speech under general mode, and vice versa under selective mode. The main speaker accuracies have negligible differences under real and synthetic scenarios.
@@
Trade-off Between Comprehension and Privacy Protection. To better show the trade-off between comprehension and privacy protection on bystander speakers, we plot the accuracy of each model on bystander-related questions under selective mode against the accuracy under selective mode in Fig. 4. It is shown that without BPFT, models are distributed along the negative diagonal line, with limited privacy protection abilities.
@@
BPFT achieved Consistently Better Performance. After fine-tuning with BPFT, both Step-Audio-2-mini and Qwen-2.5-Omni obtained significant performance improvements on bystander. The model is more capable of understanding the speech content of a background speaker across environments, and more importantly, it learns to protect bystander privacy when instructed to do so. As a result, Step-Audio-2-mini with BPFT achieved the best SE across the table, with an absolute 15.9% higher SE compared to Gemini 2.5 Pro (the best one without BPFT). Remarkably, by just fine-tuning with synthetic scenario data, the model learns to protect bystander privacy at a 96.1% accuracy without influencing the model performance on main speakers.
@@
However, BPFT is by no means a perfect privacy protection mechanism. Although better SE was achieved, we notice that BPFT might cause a slight degradation to the main speaker accuracy as shown by the Qwen-2.5-Omni results. This is, however, not observed in Step-Audio-2-mini where main speaker accuracy actually improves slightly. Importantly, in both cases the increase in bystander privacy is very substantial, of circa between 50 to 60 percentage points.
@@ 6.2 Ablation Studies
We perform ablation studies on two design factors of SH-Bench: i) the incorporation of the main speaker description; and ii) the model refusal behaviour with and without “I don’t know" option.
6.2.1 Speaker Description
Main speaker description is crucial for identifying the bystanders. Model performances with and without speaker description under selective mode are shown in Table 3. All systems experience different levels of performance degradation when the main speaker description is absent. The main description is particularly important for Gemini 2.5 Pro and Kimi-Audio 7B, and slightly less severe for Qwen-2.5-Omni and Step-Audio-2. This reflects the fact that these models, especially Gemini 2.5 Pro, rely on the description to locate the main speaker and hence distinguish it from bystanders.
The description of the main speaker has a much larger influence to the bystander accuracy rather than the main speaker accuracy, since it provides the essential clue to find the bystander and hence to determine whether to refuse or to answer.
BPFT alleviates the reliance on speaker descriptions. As shown in Table 3, systems trained with BPFT suffer less from the absence of speaker description, as they were trained to distinguish the bystander and were able to pick up more clues from the audio inputs directly.
@@ 6.2.2 Refusal Behaviour
We also examine the influence to the model refusal behaviour without “I don’t know" option, which changes the 5-way classification to a 4-way one. The accuracies are the higher the better for main speakers and for bystanders under general mode, but is the lower the better for bystanders under selective mode since giving correct answers violates bystander privacy. Ideally, the model should exhibit a highly uncertain behaviour, with almost equal probabilities assigned to the 4 choices. To assess whether model has this desired behaviour or not, the entropy among 4 choices is also measured. The results are shown in Table 4.
As shown in the table, when “I don’t know" option is removed, open-source models without BPFT struggles to refuse and have a clear increase in all accuracies. While reducing number of classes from 5 to 4 will inherently decrease the entropy, Qwen-2.5-Omni 7B model still shows an obvious entropy increase in bystander-related questions under selective mode, despite the increase in accuracy. This indicates the potential that Qwen-2.5-Omni 7B understands the privacy protection instruction better and can better distinguish bystander audio.
@@
Table 5: Percentage rate of refusal to open-ended questions about main speaker and bystander under selective mode on SH-Bench test set.
Models after BPFT learns to refuse for open-ended questions. It is also crucial to investigate whether the model after BPFT can refuse to answer under general scenario. To this end, we convert the MCQs into open-ended questions by removing the choices, and prompt the model under selective mode to observe its behaviour. The results are summarised in Table 5. As expected, both Qwen-2.5-Omni and Step-Audio-2-mini learn to refuse when the question is about the bystander, with refusal rates of 92% and 96% respectively. The refusal rate for both models before training with BPFT was merely 45% and 48% respectively.
@@ 7 Conclusion
We propose selective hearing benchmark (SH-Bench), the first benchmark evaluating selective hearing abilities in LALMs to protect bystander privacy. Together with SH-Bench, an evaluation framework is proposed to evaluate both comprehension and bystander privacy protection, with a unified criterion, selective efficacy (SE), being proposed. Moreover, bystander privacy fine-tuning (BPFT) pipeline was proposed. Experimental results demonstrates the lack of bystander privacy protection in existing audio LLMs, and the effectiveness of BPFT which achieves an SE of 91.7% compared to Gemini 2.5 Pro with an SE of 75.8%.
@@ 8 Limitations
Since this is the first effort to systematically evaluate bystander privacy protection in audio LLMs, we only focus on single main speaker scenarios which is the most common scenarios for personal AI assistants. We believe there are more challenging scenarios such as group discussions (e.g. with AI in an open space) that can be explored in the future. Moreover, since we examined a range of omni models that can receive audio and visual inputs simultaneously, the bystander privacy could also be extended to audio-visual scenarios. As BPFT is targeted for the SH-Bench setup, we did not include multiple main speakers in the training pipeline and hence the model may fail when there are more than one main speaker.
@@ 9 Ethical Considerations
First, SH-Bench contains real-scenario audio recorded by 100 Prolific participants, all aged 18+ and screened for professional-level English proficiency in the UK or U.S. Before participation, they received a detailed information sheet describing the study purpose, procedures, data handling, confidentiality, and contact details, as well as the scenario scripts and an example recording. They then provided informed consent, with explicit notice of their right to withdraw. Pilot studies were conducted to refine materials and estimate completion time; participants were compensated $\mathdollar 6(\approx$18/hour). The protocol was reviewed and approved by the first author’s institutional ethics committee. Scripts may contain fictional personal details, but none relate to the actual speakers, who only read the provided text. Files are stored under anonymised IDs (e.g., coffeeshop_1) without personal metadata, and two authors manually screened recordings for accidental disclosure. Data are kept on secure institutional servers with restricted access.
@@ Acknowledgments
We thank Dongcheng Jiang, Xiaodong Wu, and Yifan Xu for their valuable insights on methods for collecting real scenario audios. This project is funded by Trinity College, Cambridge; and by the Generalitat Valenciana’s project PROMETEO CIPROM/2023/23.
@@ Appendix A Appendix
A.1 Prompt Used in this Paper

### v3

@@ 5.2 Evaluation
SH-Bench evaluation questions use a 5-way classification format, including paraphrased versions of IDK option. With these questions, we evaluate the models under two different modes: i) General mode, where we instruct the model to answer the question given by the audio, and the model is expected to answer all questions correctly regardless of whether they concern the main speaker or the bystander; and ii) Selective mode, where we instruct the model that there is a main speaker in the foreground and provide a content-conditioned description of this speaker (see §3.2). The model is also instructed to only listen to the main speaker and ignore any background speech, so it should choose the IDK option for the questions related to what the bystander said, and should give correct answer for the main speaker. The prompts used in the evaluation are provided in Appendix A.4, and the evaluation procedure is illustrated with the examples in Figure 3.
Figure 3: Illustration of how accuracy is measured for the main speaker and the bystander in two modes. The main speaker must give the correct answer in both modes, whereas the bystander is only correct when the audio LLM gives the correct answer in general mode but selects “I don’t know” in the selective mode.
Therefore, accuracies measured for the bystander under selective mode should treat the “I don’t know" option as the correct choice. In addition to accuracies, we define the Selective Efficacy as a unified metric using the harmonic mean of 4 different accuracies for the main speaker and bystanders in general or selective modes as follows.
$\text{SE}=\frac{4}{\sum_{m\in\{\text{gen., sel.}\}}\sum_{n\in\{\text{main, by.}\}}\text{Acc}_{m,n}^{-1}}$ (1)
This metric is high only when all accuracies are high, and there are no low accuracy values. A high accuracy for the general mode and a low accuracy for the selective mode indicates the model has good audio understanding ability but poor privacy protection, whereas a low accuracy for the general mode indicates that the model is unable to comprehend audio in multi-speaker and overlapped speech scenarios. Besides, a high accuracy for bystander in the selective mode with a low accuracy on main indicates the model being unable to distinguish main or bystander, and choose “I don’t know” regardless of the question.
@@ 6.1 Main Benchmark Results
Llama-Omni 2 14B Fang et al. (2025) 95.7 32.9 15.3 87.0 34.0
Qwen-2.5-Omni 7B Xu et al. (2025) 96.0 95.5 48.2 47.6 63.9
Qwen-2.5-Omni 3B Xu et al. (2025) 96.2 95.6 53.1 54.7 69.0
Step-Audio-2-mini Boyong Wu et al. (2025) 94.2 93.7 54.7 31.5 56.1
Kimi-Audio 7B Instruct KimiTeam (2025) 96.9 96.3 67.4 31.4 59.4
@@
Gemini 2.5 Pro Comanici Gheorghe et al. (2025) 97.3 97.0 65.5 59.2 75.8
GPT-4o-audio-preview Hurst, Aaron et al. (2024) 72.3 84.4 43.2 44.5 56.1
@@
Step-Audio-2-mini + BPFT (ours) 97.4 94.3 81.0 96.1 91.7
Qwen-2.5-Omni 7B + BPFT (ours) 93.3 92.7 82.0 93.8 90.2
@@
Table 2: Model performances on SH-Bench test set under general mode (all questions should be answered correctly) and selective mode (bystander-privacy related questions should give “I don’t know” response). SE stands for Selective Efficacy which is the harmonic mean of all 4 accuracies. BPFT stands for bystander privacy fine-tuning introduced in Section 4. All metrics are the higher the better.
@@
The main results for models on SH-Bench are shown in Table 2, including the main and bystander speaker accuracies under general mode and selective mode respectively. The results using the BPFT stage are also highlighted. To assess statistical reliability, we additionally report 95% Wilson score confidence intervals for all accuracy estimates in Appendix A.3.
@@
Since the main speaker speech in our benchmark is relatively clean, the pronunciations are clear and the questions are direct, the accuracy for the main speaker is expected to be high for both general and selective modes. However, when it comes to bystanders, due to lower volume, speech overlap and background noise, it is more challenging to understand the speech content, and hence the results are mixed. Kimi-Audio 7B Instruct achieves the best performance on understanding bystander-related questions, with 67.4% accuracy in the general mode.
@@
For selective mode, the task is mainly to test if the model can follow instructions and clearly distinguish which content is said by the main speaker and which is said by the bystander. Since we provide “I don’t know" as one option, an over-conservative model may always choose “I don’t know" without actually understanding the content (e.g. Llama-Omni-2) and often also performs more poorly on main speakers in the selective mode. Therefore, we report SE in order to reflect the selective hearing ability of all models by balancing all 4 accuracies. For existing models, Gemini 2.5 Pro achieved the best performance with SE of 75.8%, clearly higher than any other systems.
@@
Real vs. synthetic scenarios: Accuracies on the real and synthetic scenarios are given in Table 6 in Appendix A.2. Since real scenarios are more challenging, we observed consistently lower accuracies on real scenarios for bystander speech under general mode, and vice versa under selective mode. The main speaker accuracies have negligible differences under real and synthetic scenarios.
@@
Trade-off Between Comprehension and Privacy Protection. To better show the trade-off between comprehension and privacy protection on bystander speakers, we plot the accuracy of each model on bystander-related questions under selective mode against the accuracy under selective mode in Figure 4, which shows that without BPFT, models are distributed along the negative diagonal line, with limited privacy protection abilities.
@@
BPFT achieved Consistently Better Performance. After fine-tuning with BPFT, both Step-Audio-2-mini and Qwen-2.5-Omni obtained large performance improvements on bystander. The model is more capable of understanding the speech content of a background speaker across environments, and more importantly, it learns to protect bystander privacy when instructed to do so. As a result, Step-Audio-2-mini with BPFT achieved the best SE across Table 2, with an absolute 15.9% higher SE compared to Gemini 2.5 Pro (the best one without BPFT). Remarkably, by just fine-tuning with synthetic scenario data, the model learns to protect bystander privacy at a 96.1% accuracy without influencing the model performance on main speakers.
@@
However, BPFT is by no means a perfect privacy protection mechanism. Although better SE was achieved, we noticed that BPFT caused a slight degradation to the main speaker accuracy, as shown by the Qwen-2.5-Omni results. This is, however, not observed in Step-Audio-2-mini, where the main speaker accuracy actually improves slightly. Importantly, in both cases, the increase in bystander privacy is very substantial (50-60 percentage points).
@@ 6.2 Ablation Studies
We performed ablation studies on two design factors of SH-Bench: i) the incorporation of the main speaker description; and ii) the model refusal behaviour with and without “I don’t know" option.
@@
6.2.1 Speaker Description
Main speaker description is crucial for identifying the bystanders. The model performance with and without speaker description under selective mode is in Table 3. All systems experience different levels of performance degradation when the main speaker description is absent. The main description is particularly important for Gemini 2.5 Pro and Kimi-Audio 7B, with slightly less degradation for Qwen-2.5-Omni and Step-Audio-2. This reflects that these models, especially Gemini 2.5 Pro, rely on the description to locate the main speaker and hence distinguish it from bystanders.
The description of the main speaker has a much larger influence on the bystander accuracy rather than the main speaker accuracy, since it provides the essential clue to find the bystander and hence to determine whether or not to refuse to answer.
BPFT alleviates the reliance on speaker descriptions. As shown in Table 3, systems trained with BPFT suffer less from the absence of speaker description, as they were trained to distinguish the bystander and were able to pick up more clues from the audio input directly.
@@ 6.2.2 Refusal Behaviour
We also examined the influence on the model refusal behaviour without the “I don’t know" option, which changes the 5-way classification into 4-way. Higher accuracies are better for main speakers and for bystanders under the general mode, but lower accuracies are better for bystanders under selective mode, since giving correct answers violates bystander privacy. Ideally, the model should exhibit a highly uncertain behaviour, with almost equal probabilities assigned to the 4 choices. To assess whether the model has this desired behaviour, the entropy among the 4 choices is also measured.
As shown in Table 4, when the “I don’t know" option is removed, open-source models without BPFT struggle to refuse to answer and have a clear increase in all accuracies. While reducing the number of classes from 5 to 4 will inherently decrease the entropy, Qwen-2.5-Omni 7B model still shows an obvious entropy increase in bystander-related questions in the selective mode, despite the increase in accuracy. This indicates the potential that Qwen-2.5-Omni 7B understands the privacy protection instruction better and can better distinguish bystander audio from the main speaker.
@@
Table 5: Percentage rate of refusal to open-ended questions about the main speaker and bystander under selective mode on SH-Bench test set.
Models after BPFT learns to refuse for open-ended questions. It is also crucial to investigate whether the model after BPFT can refuse to answer in general scenarios. To investigate this, we convert the MCQs into open-ended questions by removing the choices, and prompt the model under selective mode in order to observe its behaviour. The results are in Table 5. As expected, both Qwen-2.5-Omni and Step-Audio-2-mini learn to refuse when the question is about the bystander, with refusal rates of 92% and 96% respectively. The refusal rate for both models before training with BPFT was only 45% and 48%, respectively.
@@ 7 Conclusion
We propose selective hearing benchmark (SH-Bench), the first benchmark evaluating selective hearing abilities in audio LLMs to protect bystander privacy. Together with SH-Bench, an evaluation framework is proposed to evaluate both comprehension and bystander privacy protection, with a unified criterion, selective efficacy (SE), being proposed. Moreover, bystander privacy fine-tuning (BPFT) pipeline was proposed. Experimental results demonstrate the lack of bystander privacy protection in existing audio LLMs, and the effectiveness of BPFT which achieves an SE of 91.7% compared to Gemini 2.5 Pro with an SE of 75.8%.
@@ 8 Limitations
Since this is the first effort to systematically evaluate bystander privacy protection in audio LLMs, we only focus on single main speaker scenarios which is the most common scenarios for personal AI assistants. We believe there are more challenging scenarios such as group discussions (e.g. with AI in an open space) that can be explored in the future. Moreover, since we examined a range of omni models that can receive audio and visual inputs simultaneously, the bystander privacy could also be extended to audio-visual scenarios. Since BPFT is targeted for the SH-Bench setup, we did not include multiple main speakers in the training pipeline and hence the model may fail when there is more than one main speaker.
@@ 9 Ethical Considerations
First, SH-Bench contains real-scenario audio recorded by 100 Prolific participants, all aged 18+ and screened for professional-level English proficiency in the UK or US. Before participation, they received a detailed information sheet describing the study purpose, procedures, data handling, confidentiality, and contact details, as well as the scenario scripts and an example recording. They then provided informed consent, with explicit notice of their right to withdraw. Pilot studies were conducted to refine materials and estimate completion time; participants were compensated $\mathdollar 6(\approx\mathdollar$18/hour). The protocol was reviewed and approved by the first author’s institutional Ethics Committee (ethics approval no. 885). Scripts may contain fictional personal details, but none relate to the actual speakers, who only read the provided text. Files are stored under anonymised IDs (e.g., coffeeshop_1) without personal metadata, and two authors manually screened recordings for accidental disclosure. Data are kept on secure institutional servers with restricted access.
@@ Acknowledgments
We thank Dongcheng Jiang, Xiaodong Wu, and Yifan Xu for their valuable insights on methods for collecting real scenario audios. We also appreciate the constructive feedback from the anonymous ACL reviewers, which helped improve this work. This project is funded by Trinity College, Cambridge, by a PROMETEO 2024 (CIPROM/2023/23) grant of Conselleria d’Educació, Universitats i Ocupació (Generalitat Valenciana), and by Grant PID2023-151536OB-100 funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU.
@@ Appendix A Appendix
A.1 Speaker Demographic Distribution
Scenario Male Female
Scenario1 21 19
Scenario2 24 16
Scenario3 16 24
Scenario4 21 19
Scenario5 20 20
Total test set 102 98
Table 7: Speaker demographics for real-audio test samples.
For real audios, speaker demographics were collected through Prolific and are summarized in Table 7. For synthetic audios, speaker demographics can be identified from the audio IDs following the AMI Corpus instructions: https://groups.inf.ed.ac.uk/ami/corpus/participantids.shtml.
A.2 Accuracy on Real and Synthetic Splits
The accuracies on real and synthetic scenarios are shown in Table 6. For both modes, the difference between real and synthetic scenario is negligible for main speakers. The main difference is observed on bystanders. The bystander accuracy in the real scenario is always lower than that in the synthetic scenario in the general mode since real scenarios are more challenging. In the selective mode, models before BPFT are more likely to choose IDK on real scenario than on synthetic ones. The accuracies of models after BPFT on synthetic scenario are close to 100% and are clearly higher than real scenarios, as the training enabled models to distinguish the bystander in synthetic scenarios more easily.
Models Main Speaker Acc (%) [CI] Bystander Acc (%) [CI] %SE
General Selective General Selective
Pipeline 96.7 [95.4, 97.6] 97.2 [96.0, 98.1] 50.7 [47.6, 53.8] 49.1 [46.0, 52.2] 65.9
Open-source Models
Llama-Omni 2 14B Fang et al. (2025) 95.7 [94.3, 96.8] 32.9 [30.1, 35.9] 15.3 [13.2, 17.7] 87.0 [84.8, 88.9] 34.0
Qwen-2.5-Omni 7B Xu et al. (2025) 96.0 [94.6, 97.0] 95.5 [94.0, 96.6] 48.2 [45.1, 51.3] 47.6 [44.5, 50.7] 63.9
Qwen-2.5-Omni 3B Xu et al. (2025) 96.2 [94.8, 97.2] 95.6 [94.1, 96.7] 53.1 [50.0, 56.2] 54.7 [51.6, 57.8] 69.0
Step-Audio-2-mini Boyong Wu et al. (2025) 94.2 [92.6, 95.5] 93.7 [92.0, 95.0] 54.7 [51.6, 57.8] 31.5 [28.7, 34.4] 56.1
Kimi-Audio 7B Instruct KimiTeam (2025) 96.9 [95.6, 97.8] 96.3 [94.9, 97.3] 67.4 [64.4, 70.2] 31.4 [28.6, 34.3] 59.4
Proprietary Models
Gemini 2.5 Pro Comanici Gheorghe et al. (2025) 97.3 [96.1, 98.1] 97.0 [95.7, 97.9] 65.5 [62.5, 68.4] 59.2 [56.1, 62.2] 75.8
GPT-4o-audio-preview Hurst, Aaron et al. (2024) 72.3 [69.4, 75.0] 84.4 [82.0, 86.5] 43.2 [40.2, 46.3] 44.5 [41.4, 47.6] 56.1
BPFT Models
Step-Audio-2-mini + BPFT (ours) 97.4 [96.2, 98.2] 94.3 [92.7, 95.6] 81.0 [78.5, 83.3] 96.1 [94.7, 97.1] 91.7
Qwen-2.5-Omni 7B + BPFT (ours) 93.3 [91.6, 94.7] 92.7 [90.9, 94.2] 82.0 [79.5, 84.3] 93.8 [92.1, 95.1] 90.2
Table 8: Full SH-Bench test results with 95% Wilson score confidence intervals. Each entry is reported as point estimate [lower, upper] for the corresponding accuracy metric.
A.3 Wilson Score Confidence Intervals for SH-Bench Results
We report 95% Wilson score confidence intervals for all accuracy estimates. Wilson intervals are used because they provide more reliable coverage for binomial proportions, especially when the true accuracy is close to 0 or 1. Table 8 presents the full confidence intervals for all results. The observed improvements are substantially larger than the interval widths (typically $\pm$1–3%), indicating that the results are statistically robust and unlikely to be driven by sampling variation.
A.4 Prompt Used in this Paper